ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

3. Number Theory and Arithmetic Algebraic Geometry

1991 MS Classification: not given Bolger, Robert E., Fairfield University, Connecticut, USA Fermat’s last theorem: Was his proof marginal The idea behind his proof for n = 3, 4! On the other hand, it is very hard to believe that his proof was valid geometrically!